OpenStudy (anonymous):

HELP PLEASE Sara charges $2.25 per an hour when she babysits plus $5.00 for transportation expenses. Which function rule represents the amount of y Sara charges to babysit for X hours?

OpenStudy (anonymous):

2.25h+5.00?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Options are A- y=2.25 + 5.00 B- y=5.00y + 2.25 C- y=5.00x + 2.25 D- y= 7.25x

OpenStudy (anonymous):

y = 5.00 + 2.25 x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Your welcome.